# Parallel transport on matrix manifolds and Exponential Action

### v1

@@ 9. Numerical studies
As mentioned, the algorithm [2] for exponential action requires a norm estimate for the operator $\mathrm{P}$. For Stiefel and flag manifolds, after a rescaling, this operator becomes antisymmetric in the Frobenius norm. Numerically, in this case, the estimate using just the 1-norm $\|\mathrm{P}\|_{1}$ is adequate, instead of a higher $p$-norm estimate.
@@
Our next observation is for our problem, $\|\mathrm{P}\|_{1}$ can be estimated more efficiently than by calling a library function. Recall that on a $n$-dimension vector space with a basis $\{e_{i}\}_{i=1}^{n}$, the 1-norm of a vector $\sum c_{i}e_{i}$ is $\sum_{i}|c_{i}|$. Vectorizing an operator $\mathrm{P}$ using this basis, we have $\|\mathrm{P}\|_{1}=\max_{j}|\mathrm{P}e_{j}|_{1}$, which has a cost of $O(n^{2})$, thus, it is $O(m^{2}d^{2})$ on a matrix space $\mathbb{R}^{m\times d}$. The sampling in [10] reduces this cost, but we can reduce it further by exploiting the structure of $\mathrm{P}$. For the Stiefel and flag manifold cases, we provide an estimate based on norms of $A$ and $R$. The resulting customized implementation of exponential action could be more than $10$ times faster than the library version, significantly reducing the total time cost of the parallel transport calculation. To keep focus, we do not attempt these estimates of $\|\mathrm{P}\|_{1}$ for $\operatorname{GL}^{+}(n)$ and $\mathrm{SO}(n)$, while they seem likely to be available. Likewise, we will not discuss the scalings to make $\mathrm{P}$ antisymmetric for these manifolds as in lemma 4 below.
@@
Further, the scipy function $expm{\_}multiply$ is not yet implemented in JAX [4], a Python library that provides easy compilation to run in GPU and allows automatic differentiation (AD). We use our estimates to implement a simple version of $expv$ in JAX, and use the AD feature to compute the derivative of the parallel transport, verifying that our implementation is highly accurate. We offer $expm{\_}multiply$-based parallel transport for all four manifolds while providing customized implementations for Stiefel and flag manifolds, with AD(JAX) and without-AD versions. The discussion here will likely apply to other languages using similar algorithms.
@@
First, for a real number $c$, define $s_{c}$ to be the operator on $\mathbb{R}^{m\times d}$ with $m\geq d$, scaling the top $d\times d$ block by $c$ and preserving the bottom $(m-d)\times d$ block. We have $\operatorname{expm}\mathrm{P}=s_{c}^{-1}\operatorname{expm}(s_{c}\mathrm{P}s_{c}^{-1})s_{c}$ for an operator $\mathrm{P}$ on $\mathbb{R}^{m\times d}$. The following computes an upper bound for $\|s_{\alpha^{\frac{1}{2}}}\mathrm{P}_{AR}s_{\alpha^{-\frac{1}{2}}}\|_{1}$ with a cost of $O(d^{2})$:
@@
For $k\geq 0$ and $A=(a_{ij})_{ij=(1,1)}^{(d,d)}\in\mathfrak{so}(d),R=(r_{ij})_{ij=(1,1)}^{(k,d)}\in\mathbb{R}^{k\times d}$, with respect to the standard basis of $\mathbb{R}^{(d+k)\times d}$, the operator $\mathrm{P}_{AR}^{bal}:=s_{\alpha^{\frac{1}{2}}}\circ\mathrm{P}_{AR}\circ s_{\alpha^{-\frac{1}{2}}}$ with $\mathrm{P}_{AR}$ in eq. 52 operating on $\mathbb{R}^{(d+k)\times d}$ has an upper bound on the $1$-norm
(56) $\begin{gathered}\|\mathrm{P}_{AR}^{bal}\|_{1}=\max_{|w|_{1}=1}\lvert\mathrm{P}_{AR}^{bal}w\rvert_{1}\leq\max(n_{A},n_{R})\\ \text{where }n_{A}=\max_{j}(\alpha^{\frac{1}{2}}\sum_{i=1}^{d}(\lvert r_{ij}\rvert+\lvert 4\alpha-1\rvert\|A\|_{1})),\\ n_{R}=\max_{j}(\alpha\sum_{i}(\lvert a_{ij}\rvert+\alpha^{\frac{1}{2}}\|R\|_{\infty})).\end{gathered}$
Also, define the operator $\mathsf{skew}_{top}$ on $\mathbb{R}^{(d+k)\times d}$ to be the map antisymmetrizing the top $d\times d$ block, then $\mathrm{P}_{AR}^{bal}\circ\mathsf{skew}_{top}:\begin{bmatrix}w_{a}\\ w_{r}\end{bmatrix}\mapsto\mathrm{P}_{AR}^{bal}\begin{bmatrix}(w_{a})_{\mathrm{skew}}\\ w_{r}\end{bmatrix}$ is antisymmetric in the Frobenius inner product. With $\mathrm{P}^{\operatorname{Flag}}_{AR}$ given in eq. 55 at $\alpha=\frac{1}{2}$, we have the same bound for $s_{\alpha^{\frac{1}{2}}}\circ\mathrm{P}^{\operatorname{Flag}}_{AR}\circ s_{\alpha^{-\frac{1}{2}}}$.
@@
For $w=\begin{bmatrix}w_{a}\\ w_{r}\end{bmatrix}$, we have
$\mathrm{P}_{AR}^{bal}w=\begin{bmatrix}\alpha^{\frac{1}{2}}((4\alpha-1)\alpha^{-\frac{1}{2}}w_{a}A+R^{\mathsf{T}}w_{r})_{\mathrm{skew}}\\ \alpha(w_{r}A-\alpha^{-\frac{1}{2}}Rw_{a})\end{bmatrix}=\begin{bmatrix}((4\alpha-1)w_{a}A+\alpha^{\frac{1}{2}}R^{\mathsf{T}}w_{r})_{\mathrm{skew}}\\ \alpha w_{r}A-\alpha^{\frac{1}{2}}Rw_{a}\end{bmatrix}.$
@@
where |.$|.|_{\vec{1}}$ denotes the $1$-norm after vectorization, because $E_{ij}E_{jr}=E_{ir}$ is distinct from $E_{sj}E_{ji}=E_{si}$ unless $i=r=s$, in this case, we have a cancelation. For $i>d$
@@
This gives us the bound. To see $\mathrm{P}_{AR}^{bal}\circ\mathsf{skew}_{top}$ is antisymmetric, note the diagonal parts $w_{a}\mapsto(4\alpha-1)((w_{a})_{\mathrm{skew}}A)_{\mathrm{skew}}$ and $w_{r}\mapsto\alpha w_{r}A$ are both antisymmetric, since $A$ is antisymmetric. We can verify directly the off-diagonal operators $F_{1}:w_{r}\mapsto\alpha^{\frac{1}{2}}(R^{\mathsf{T}}w_{r})_{\mathrm{skew}}$ and $F_{2}:w_{a}\mapsto-\alpha^{\frac{1}{2}}R(w_{a})_{\mathrm{skew}}$ satisfy $F_{1}=-F_{2}^{\mathsf{T}}$. For flag manifolds, clearing flag diagonal blocks does not increase norms, thus the same estimate holds. ∎
@@
With this estimate, following scipy’s implementation, we have a finer extension of [2, table 3.1] containing a look up table for constants $\theta_{m}$ from $m$. The number of Taylor terms $m_{*}$ will be the value of $m$ minimizing $\mathrm{ceil}(\frac{m\|\mathrm{P}\|_{1}}{\theta_{m}})$ and the scaling parameter $s$ is set to be $\mathrm{ceil}(\frac{\|\mathrm{P}\|_{1}}{\theta_{m_{*}}})$, both are used in the approximation $\operatorname{expv}(\mathrm{P},\mathtt{b})\approx(T_{m}(s^{-1}\mathrm{P}))^{s}\mathtt{b}$.
In table 2, we show the computational time along a time grid, for different $n$ and $\alpha$ for Stiefel manifolds, and for $\alpha=\frac{1}{2}$ (the canonical metric). The flag manifold implementation simplifies the transport formula when $\alpha=\frac{1}{2}$, resulting in slightly faster execution. Table table 3 shows execution time by $d$, also along the same time grid. Using automatic differentiation, in [22] we can verify that these correspond to highly accurate computation, depending on the manifold, with the transport equation satisfied with accuracy around $10^{-10}$ to $10^{-12}$.
Stiefel man. comp. time (s) Flag man. comp. time.(s)
Time grid t(s) for geodesics Time grid t(s) for geodesics
$\alpha$ n 0.5 1.0 2.0 5.0 20.0 0.5 1.0 2.0 5.0 20.0
$\frac{1}{2}$ 100 0.04 0.08 0.14 0.25 0.59 0.04 0.07 0.12 0.17 0.50
200 0.07 0.13 0.16 0.25 0.75 0.06 0.10 0.16 0.23 0.56
1000 0.13 0.19 0.25 0.48 1.23 0.11 0.17 0.21 0.34 1.33
2000 0.23 0.28 0.35 0.61 1.80 0.23 0.27 0.32 0.54 1.60
1 100 0.08 0.12 0.17 0.32 1.13
200 0.09 0.14 0.20 0.37 1.44
1000 0.17 0.21 0.33 0.61 2.10
2000 0.26 0.32 0.46 0.76 2.73
Table 2. Computational time for parallel transport for Stiefel and flag manifolds by $n$ for $d=50$. The flags are $(20,15,15,n-d)$. The canonical metric corresponds to $\alpha=\frac{1}{2}$, the Euclidean embedded metric corresponds to $\alpha=1$. The table shows approximate linear time in $n$ and $t$. Execution is on a free colab session.
Stiefel man. comp. time (s) Flag man. comp. time.(s)
Geo. time grid t(s) Geo. time grid t(s)
$\alpha$ d 0.5 1.0 2.0 5.0 20.0 0.5 1.0 2.0 5.0 20.0
$\frac{1}{2}$ 5 0.03 0.04 0.06 0.13 0.29
10 0.03 0.05 0.08 0.15 0.40 0.03 0.05 0.10 0.15 0.37
20 0.06 0.08 0.14 0.19 0.45 0.06 0.08 0.12 0.17 0.41
100 0.35 0.59 0.97 2.15 8.39 0.30 0.48 0.70 2.00 7.19
1 5 0.04 0.07 0.11 0.15 0.40
10 0.04 0.06 0.09 0.17 0.61
20 0.07 0.10 0.16 0.26 0.66
100 0.54 0.76 1.54 3.51 14.07
Table 3. Computational time for parallel transport for Stiefel and flag manifolds by $d$ for $n=1000$. We scale the partial flag of $(d_{1},d_{2},d_{3})=(4,3,3)$ by $\frac{d}{4+3+3}$ to $d$ for flag manifolds, setting $d_{4}=n-d$. Computational time is nonlinear in $d$.
9.1. Parallel transport as isometries
We also demonstrate numerically the isometric property of parallel transport as implemented by our method, by showing the inner product matrix of a set of vectors is preserved during transportation. We consider a geodesic starting at $\operatorname{I}_{2000,100}$, with twenty tangent vectors having integer lengths sampled from integers between $1$ and $60$, while the relative angles are randomly generated. The geodesic initial velocity vector $v$ has length $1$. We compute the inner product matrix of the random tangent vectors at $t=0$ and of the transported vectors along a time grid. We measure the maximum absolute difference between the inner product matrices at each grid point and the initial point and plot the $\log_{10}$ of the differences in fig. 1. The absolute errors are around $10^{-12}$ up to $t=15$.
(a) Stiefel manifolds
(b) Flag manifolds
[…]
10. Discussion
In this paper, we provide an efficient algorithm to compute parallel transport on Stiefel and flag manifolds, as well as transport formulas for the generalized linear group and the special orthogonal group. We also provide a framework for parallel transports on groups with transposable Lie algebra. The symplectic group [6] also has transposable Lie algebra, thus the framework here should apply. We believe the framework developed here will be useful for other manifolds in applied mathematics.
Acknowledgments
Du Nguyen would like to thank his family for their support in this project. Stefan Sommer is supported by a research grant (VIL40582) from VILLUM FONDEN.
Appendix A The Grassmann case

### v2

@@ Appendix A Exponential action
In these implementations, the $1$-norm is estimated using the algorithm in [14]. We will use expm_multiplyexpm{\_}multiply from scipy for $\operatorname{GL}(n)$ and $\mathrm{SO}(n)$, while we will show a customized implementation based on efficient estimates of the $1$-norm of the operator $\mathrm{P}$ for Stiefel and flag manifolds in lemma 6 could speed up the algorithm significantly, without the higher $p$ estimates. In particular, we show for these manifolds, after rescaling, the operator $\mathrm{P}$ becomes antisymmetric in the Frobenius norm. Numerically, in this case, the estimate using just the 1-norm $\|\mathrm{P}\|_{1}$ is adequate.
@@
Our next observation is for our problem, $\|\mathrm{P}\|_{1}$ can be estimated more efficiently than by calling a library function. Recall that on a $n$-dimension vector space with a basis $\{e_{i}\}_{i=1}^{n}$, the 1-norm of a vector $\sum c_{i}e_{i}$ is $\sum_{i}|c_{i}|$. Vectorizing an operator $\mathrm{P}$ using this basis, the operator $1$-norm is given by $\|\mathrm{P}\|_{1}=\max_{j}|\mathrm{P}e_{j}|_{1}$, which has a cost of $O(n^{2})$, thus, it is $O(m^{2}d^{2})$ on a matrix space $\mathbb{R}^{m\times d}$. The sampling in [14] reduces this cost, but we can reduce it further by exploiting the structure of $\mathrm{P}$. For the Stiefel and flag manifold cases, we provide an estimate based on norms of $A$ and $R$. The resulting customized implementation of exponential action could be more than $10$ times faster than the library version, significantly reducing the total time cost of the parallel transport calculation. To keep focus, we do not attempt these estimates of $\|\mathrm{P}\|_{1}$ for $\operatorname{GL}^{+}(n)$ and $\mathrm{SO}(n)$ (the latter uses a similar estimate to Stiefel manifolds). Likewise, we will not discuss the scalings to make $\mathrm{P}$ antisymmetric in the Frobenius norm for these manifolds as in lemma 6 below.
@@
Further, the scipy function expm_multiplyexpm{\_}multiply is not yet implemented in JAX [4], a Python library that provides easy compilation to run in GPU and allows automatic differentiation (AD). We use our estimates to implement a simple version of expvexpv in JAX, and use the AD feature to compute the derivative of the parallel transport, verifying that our implementation is highly accurate. We offer expm_multiplyexpm{\_}multiply-based parallel transport for all four manifolds while providing customized implementations for Stiefel and flag manifolds, with AD(JAX) and without-AD versions. The discussion here will likely apply to other languages using similar algorithms.
@@
First, for a real number $c$, define $\mathsf{S}_{c}$ to be the operator on $\mathbb{R}^{m\times d}$ with $m\geq d$, scaling the top $d\times d$ block by $c$ and preserving the bottom $(m-d)\times d$ block. We have $\operatorname{expm}\mathrm{P}=\mathsf{S}_{c}^{-1}\operatorname{expm}(\mathsf{S}_{c}\mathrm{P}\mathsf{S}_{c}^{-1})\mathsf{S}_{c}$ for an operator $\mathrm{P}$ on $\mathbb{R}^{m\times d}$. The following computes an upper bound for $\|\mathsf{S}_{\alpha^{\frac{1}{2}}}\mathrm{P}_{AR}\mathsf{S}_{\alpha^{-\frac{1}{2}}}\|_{1}$ with a cost of $O(d^{2})$:
@@
For $k\geq 0$ and $A=(a_{ij})_{ij=(1,1)}^{(d,d)}\in\mathfrak{so}(d),R=(r_{ij})_{ij=(1,1)}^{(k,d)}\in\mathbb{R}^{k\times d}$, with respect to the standard basis of $\mathbb{R}^{(d+k)\times d}$, the operator $\mathrm{P}_{AR}^{bal}:=\mathsf{S}_{\alpha^{\frac{1}{2}}}\circ\mathrm{P}_{AR}\circ\mathsf{S}_{\alpha^{-\frac{1}{2}}}$ with $\mathrm{P}_{AR}$ in eq. 62 operating on $\mathbb{R}^{(d+k)\times d}$ has an upper bound on the $1$-norm
(66) $\begin{gathered}\|\mathrm{P}_{AR}^{bal}\|_{1}=\max_{|\mathtt{W}|_{1}=1}\lvert\mathrm{P}_{AR}^{bal}\mathtt{W}\rvert_{1}\leq\max(n_{A},n_{R})\\ \text{where }n_{A}=\max_{j}(\alpha^{\frac{1}{2}}\sum_{i=1}^{d}(\lvert r_{ij}\rvert+\lvert 4\alpha-1\rvert\|A\|_{1})),\\ n_{R}=\max_{j}(\alpha\sum_{i}(\lvert a_{ij}\rvert+\alpha^{\frac{1}{2}}\|R\|_{\infty})).\end{gathered}$
Also, define the operator $\mathsf{skew}_{top}$ on $\mathbb{R}^{(d+k)\times d}$ to be the map antisymmetrizing the top $d\times d$ block, then $\mathrm{P}_{AR}^{bal}\circ\mathsf{skew}_{top}:\begin{bmatrix}\mathtt{W}_{a}\\ \mathtt{W}_{r}\end{bmatrix}\mapsto\mathrm{P}_{AR}^{bal}\begin{bmatrix}(\mathtt{W}_{a})_{\mathrm{skew}}\\ \mathtt{W}_{r}\end{bmatrix}$ is antisymmetric in the Frobenius inner product. With $\mathrm{P}^{\operatorname{Flag}}_{AR}$ given in eq. 65 at $\alpha=\frac{1}{2}$, we have the same bound for $\mathsf{S}_{\alpha^{\frac{1}{2}}}\circ\mathrm{P}^{\operatorname{Flag}}_{AR}\circ\mathsf{S}_{\alpha^{-\frac{1}{2}}}$.
@@
For $\mathtt{W}=\begin{bmatrix}\mathtt{W}_{a}\\ \mathtt{W}_{r}\end{bmatrix}$, we have
$\mathrm{P}_{AR}^{bal}\mathtt{W}=\begin{bmatrix}\alpha^{\frac{1}{2}}((4\alpha-1)\alpha^{-\frac{1}{2}}\mathtt{W}_{a}A+R^{\mathsf{T}}\mathtt{W}_{r})_{\mathrm{skew}}\\ \alpha(\mathtt{W}_{r}A-\alpha^{-\frac{1}{2}}R\mathtt{W}_{a})\end{bmatrix}=\begin{bmatrix}((4\alpha-1)\mathtt{W}_{a}A+\alpha^{\frac{1}{2}}R^{\mathsf{T}}\mathtt{W}_{r})_{\mathrm{skew}}\\ \alpha\mathtt{W}_{r}A-\alpha^{\frac{1}{2}}R\mathtt{W}_{a}\end{bmatrix}$.
@@
where $|.|_{\vec{1}}$ denotes the $1$-norm after vectorization, because $E_{ij}E_{jr}=E_{ir}$ is distinct from $E_{sj}E_{ji}=E_{si}$ unless $i=r=s$, in this case, we have a cancellation. For $i>d$
@@
This gives us the bound. To see $\mathrm{P}_{AR}^{bal}\circ\mathsf{skew}_{top}$ is antisymmetric, note the diagonal parts $\mathtt{W}_{a}\mapsto(4\alpha-1)((\mathtt{W}_{a})_{\mathrm{skew}}A)_{\mathrm{skew}}$ and $\mathtt{W}_{r}\mapsto\alpha\mathtt{W}_{r}A$ are both antisymmetric in the Frobenius norm, since $A$ is antisymmetric. We can verify directly the off-diagonal operators $F_{1}:\mathtt{W}_{r}\mapsto\alpha^{\frac{1}{2}}(R^{\mathsf{T}}\mathtt{W}_{r})_{\mathrm{skew}}$ and $F_{2}:\mathtt{W}_{a}\mapsto-\alpha^{\frac{1}{2}}R(\mathtt{W}_{a})_{\mathrm{skew}}$ satisfy $F_{1}=-F_{2}^{\mathsf{T}}$ (the operator transpose is also in the Frobenius norm). For flag manifolds, clearing flag diagonal blocks does not increase norms, thus the same estimate holds. ∎
@@
With this estimate, following scipy’s implementation, we have a finer extension of [2, table 3.1] containing a look up table for constants $\theta_{m}$ from $m$. The number of Taylor terms $m_{*}$ will be the value of $m$ minimizing $\mathrm{ceil}(\frac{m\|\mathrm{P}\|_{1}}{\theta_{m}})$ and the scaling parameter $s$ is set to be $\mathrm{ceil}(\frac{\|\mathrm{P}\|_{1}}{\theta_{m_{*}}})$, both are used in the approximation $\operatorname{expv}(\mathrm{P},\mathtt{B})\approx(T_{m}(s^{-1}\mathrm{P}))^{s}\mathtt{B}$. In algorithm 2 We implement an evaluation of $\operatorname{expv}$, given an estimate of the number of Taylor terms and the scaling factor using the 1-norm estimate, $\mathrm{P}$ is assumed to be an antisymmetric operator in the Frobenius norm (if this does not hold, convergence may be slower than the library version which uses higher norms.)
Algorithm 2 Computing exponential action $e^{t\mathrm{P}}B$ based on the scaling method. The number of Taylor terms $m_{*}$ and the scaling factor $s_{*}$ are pre-estimated.
1:Input: $\mathrm{P},B,t,m_{*},s_{*}$
2:Choose a relative tolerance level toltol
3:for $i=0,\cdots,s_{*}-1$ do
4:  $old\_norm\leftarrow|B|_{2}$
5:  $F\leftarrow B$
6:  for $j=0,\cdots,m_{*}-1$ do
7:   $B\leftarrow\frac{t}{s_{*}(j+1)}\mathrm{P}(B)$
8:   $new\_norm=|B|_{2}$
9:   $F\leftarrow F+B$
10:   if $old\_norm+new\_norm<tol|F|_{2}$ then
11:     break
12:   end if
13:   old_norm=new_normold\_norm=new\_norm
14:  end for
15:end for
If we use the library expm_multiply function, we need to supply the transpose operator for $\mathrm{P}_{\mathtt{V}}$ in eq. 29. We have
Lemma 7.
If $\mathfrak{a}$ is a transposable subalgebra and $\mathcal{B}=\beta\beta_{0}\operatorname{I}_{\mathfrak{a}}$, in the Frobenius inner product, for $\mathtt{C}\in\mathcal{E}=\mathbb{R}^{n\times n}$, the adjoint of the operator $\mathrm{P}_{\mathtt{V}}$ in eq. 29 is given by
(67) $\mathrm{P}_{\mathtt{V}}^{\mathsf{T}}\mathtt{C}=\frac{1}{2}\left([\mathtt{C},\mathtt{V}^{\mathsf{T}}]+(1+\beta)([(\mathcal{P}_{\mathfrak{a}}\mathtt{V})^{\mathsf{T}},\mathtt{C}]-\mathcal{P}_{\mathfrak{a}}[\mathtt{C},\mathtt{V}^{\mathsf{T}}])\right).$
Proof.
We use the equality $\operatorname{Tr}\mathtt{B}_{1}^{\mathsf{T}}[\mathtt{B}_{2},\mathtt{X}]=\operatorname{Tr}\mathtt{B}_{2}^{\mathsf{T}}[\mathtt{B}_{1},\mathtt{X}^{\mathsf{T}}]$ for $\mathtt{B}_{1},\mathtt{B}_{2},\mathtt{X}\in\mathbb{R}^{n\times n}$ by taking transpose then use eq. 10. Thus, for $\mathtt{B}\in\mathcal{E}=\mathbb{R}^{n\times n}$,
$\begin{gathered}\operatorname{Tr}\mathtt{C}^{\mathsf{T}}\left(\frac{1}{2}[\mathtt{B},\mathtt{V}]+\frac{1+\beta}{2}\left(-[\mathcal{P}_{\mathfrak{a}}\mathtt{B},\mathtt{V}]-[\mathtt{B},\mathcal{P}_{\mathfrak{a}}\mathtt{V}]\right)\right)\\ =\frac{1}{2}\operatorname{Tr}\mathtt{B}^{\mathsf{T}}[\mathtt{C},\mathtt{V}^{\mathsf{T}}]-\frac{1+\beta}{2}\left(\operatorname{Tr}(\mathcal{P}_{\mathfrak{a}}\mathtt{B})^{\mathsf{T}}[\mathtt{C},\mathtt{V}^{\mathsf{T}}]+\operatorname{Tr}\mathtt{B}^{\mathsf{T}}[\mathtt{C},(\mathcal{P}_{\mathfrak{a}}\mathtt{V})^{\mathsf{T}}]\right).\end{gathered}$
Using $\operatorname{Tr}(\mathcal{P}_{\mathfrak{a}}\mathtt{B})^{\mathsf{T}}[\mathtt{C},\mathtt{V}^{\mathsf{T}}]=\operatorname{Tr}\mathtt{B}^{\mathsf{T}}(\mathcal{P}_{\mathfrak{a}}[\mathtt{C},\mathtt{V}^{\mathsf{T}}])$, we get the expression for $\mathrm{P}^{\mathsf{T}}_{\mathtt{V}}$. ∎
Appendix B The Grassmann case